# Voting-Based Semi-Parallel Proof-of-Work Protocol

### v3

@@ Abstract
Parallel Proof-of-Work (PoW) protocols are suggested to improve the safety guarantees, transaction throughput and confirmation latencies of Nakamoto consensus. In this work, we first consider the existing parallel PoW protocols and develop hard-coded incentive attack structures. Our theoretical results and simulations show that the existing parallel PoW protocols are more vulnerable to incentive attacks than the Nakamoto consensus, e.g., attacks have smaller profitability threshold and they result in higher relative rewards. Next, we introduce a voting-based semi-parallel PoW protocol that outperforms both Nakamoto consensus and the existing parallel PoW protocols from most practical perspectives such as communication overheads, throughput, transaction conflicts, incentive compatibility of the protocol as well as a fair distribution of transaction fees among the voters and the leaders. We use state-of-the-art analysis to evaluate the consistency of the protocol and consider Markov decision process (MDP) models to substantiate our claims about the resilience of our protocol against incentive attacks.
I Introduction
@@
We introduce several rules for honest miners to make sure picking the proposal ledger that offers the most fees becomes a Nash equilibrium for all miners of size $\alpha<0.5$ as $L$ increases. As a result, only one hidden ledger is shared in the equilibrium and communication overheads as well as transaction conflicts are avoided. We introduce punishment mechanisms as in [20, 18, 19] to further reinforce incentive compatibility of the protocol as well as a fair distribution of transaction fees among the voters and the leaders. The introduction of ledger choice rule and split rule for the distribution of transaction fees makes sure that arbitrary manipulations of ledgers are not feasible for adversaries. We adapt the state-of-the-art security latency analysis of Nakamoto consensus [31, 32] to analyze the double-spending probability for our protocol. Further, we consider Markov decision process (MDP) models to substantiate our claims about the resilience of our protocol against incentive attacks.
In the rest of this paper, we assume a single colluding adversarial entity which holds $\alpha$ fraction of the hashpower in the system. We assume a network of $n$ honest miners connected to each other with some network topology and large $n$. We denote the network influence of the adversary, i.e., the fraction of honest miners that prefer the adversarial chain in case of a fork as $\gamma$. We abstract the mining puzzles with a collision-free hash function, $f_{H}$, which acts as a random oracle. In Section II, we summarize the existing variations of the parallel PoW protocols and show their vulnerabilities against incentive attacks. In Section III, we formally introduce our protocol and subsequently analyze it in Section IV. In Section V, we discuss how our protocol relates to the existing work on PoW based protocols in general, its limitations and possible future work.
II Incentive Attacks on Existing Variations of Parallel PoW Protocols
@@
II-A Bobtail
In Bobtail, each proof has a hash value and can be created in parallel as usual and any proof with a hash value lower than a certain threshold is a valid proof. [^1] In the actual paper of [4], the average hash value of $L$ proofs needs to be below a threshold, however, for practical purposes and simplicity, we ignore proofs above a certain threshold and assume any proof below it is a valid proof as in [33]. The proof that has the lowest hash value is allowed to combine the current $L$ proofs to create a block. To prevent the proof withholding attacks, each proof has a support value, which is the main difference to the raw form. This support value is essentially a pointer to the lowest hash valued proof it has seen so far among the proofs for the current block height. To create a block, a lowest hash proof is only allowed to combine the proofs that have higher hash and support hash value.
@@ II-A Bobtail
To further strengthen the protocol, the authors consider a bonus reward rule where proofs, whose support proof is the lowest hash valued proof which eventually created the final block, get a bonus reward. The protocol also defines further incentive rules about prioritizing inclusion of proofs in case multiple sets of $L$ proofs are possible. We refer the reader to [4] for further details.
In Appendix A, we first provide details of the two attacks considered in the original paper of Bobtail, and later create a stronger hard-coded attack for which a pseudo-code is also given. We also provide both a theoretical overview as well as a simulation result which shows Bobtail simply has profitability threshold of $\alpha_{T}=0$ due to the informational advantage that the adversary possesses. As a result, even a small home miner can mount an incentive attack on the protocol. It is also noted that the attack scales block interval times, decreases the throughput and increases the latency.
II-B Tree- and DAG-Style Voting
Tailstorm [29] is a tree-style parallel PoW protocol [30]. Tailstorm introduces two main innovations compared to the raw form of the parallel PoW protocol. First, the proofs that are being generated for the current block height can refer to each other, creating a tree of proofs. As long as the root of the tree is the previous block, they are all eligible to be combined together to create the next block. Tailstorm further deviates from the raw form in how it rewards the proofs. Each block is rewarded proportional to the maximal depth of its proof tree and the reward is distributed equally to each proof in the tree. Note that in Tailstorm, every proof in a block gets rewarded equally, which also means that they are punished equally, since they can get more rewards if they build their proofs in a sequential chain and if a proof creates a fork, the reward for every proof inside the block decreases equally.
The addition of this tree structure and the new reward mechanism aim to mitigate the proof withholding attacks. DAG-style voting [30] introduces modifications along the same lines that aim to fix the unfair punishment mechanism in Tailstorm. In DAG-style voting, proofs are allowed to refer to multiple proofs which results in DAG. In this case, the reward of each proof is proportional to the total number of ancestor and descendant proofs it has within the block, which punishes the non-sequentiality more fairly. We refer the reader to [30] for further details as needed.
@@ II-B Tree- and DAG-Style Voting
II-C Numerical Results for the Incentive Attacks
@@
The handcrafted attacks we consider work irrespective of $\gamma$ value, hence better attacks exist that take advantage of $\gamma$. Moreover, the handcrafted attacks are non-dynamic in the sense that the adversary does not change its strategy even when it encounters a high variance mining time. For example, in selfish mining attack of Nakamoto consensus, the profit of the attack comes from the fact that the adversary takes advantage of the high variance of the mining times, i.e., whenever the honest miners encounter a high block interarrival time, the adversary has an opportunity to be conversely lucky [4]. For example, since rewards are scaled according to the depth, the adversary could benefit from dynamically switching to honest mining for blocks for which it has created less than usual proofs.
III Proposed Protocol Description
@@
Notice that, if a protocol separates the proofs from the ledgers, it decreases propagation times of proofs. As a result, less proofs will be mined during delay periods and thus the level of parallelness does not need to be high. This however, requires a ledger election mechanism that is incentive compatible. Further, in this case, ledger conflicts can be omitted with a clever mechanism as well. In this section, we explain a protocol that allows a reasonable amount of parallelness with small proof propagation times, an incentive compatible leader election mechanism, less communication overhead and has no ledger conflicts. Further, it has the superior variance properties compared to Nakamoto consensus, i.e., a more fair distribution of coins, better safety guarantees and resilience against withholding attacks and throughput increase.
III-A Blocks and Chain Structure
@@
A valid initiator proof consists of 6 parts concatenated together, i.e., $\omega^{h}_{1}=(\omega^{h}_{1,1},\ldots,\omega^{h}_{1,6})$ where w.l.o.g. we assume the miner who mines $\omega^{h}_{1}$ is the miner $i=1$. The sixth part of the proof, i.e., $\omega^{h}_{1,6}$, contains the nonce to make sure $f_{H}(\omega^{h}_{1})<T_{L}$. The fifth part, i.e., $\omega^{h}_{1,5}$, is simply the hash of the coin address of the miner $i=1$ to store its mining rewards in case its solution is accepted. The fourth part, $\omega^{h}_{1,4}$, is the Merkle root of the transaction ledger proposed by miner $1$, which will be explained in detail later. The third part, $\omega^{h}_{1,3}$, is simply the total accumulated transaction fees of the proposed ledger associated with $\omega^{h}_{1,4}$. The second part is a summary of (w.l.o.g. assume $L$) clustered proofs of preceding block height, i.e., $\omega^{h}_{1,2}=\omega^{h-1}=(f_{H}(\omega^{h-1}_{1}),\ldots,f_{H}(\omega^{h-1}_{L}))$. If two initiator proofs at height h1h-1 exist, say miner $1$ and miner $2$ created valid initiator proofs $\omega^{h}_{1}$ and $\omega^{h}_{2}$, then, they need to agree about the history to be clustered together in a block in order to prevent double-spending, i.e., $\omega^{h-1}_{1,2}=\omega^{h-1}_{2,2}$ ($\omega^{h-1}_{1,1}=\omega^{h-1}_{2,1}$ should also be satisfied, which will be explained in Section III-B).
@@
III-A1 Example proof model
@@
In a similar manner, at block height h1h-1, the proof represented with a red circle cannot be combined with the proof represented with a blue circle, since their parent sets differ. Regarding the red circles of block height $h+2$, multiple scenarios are possible. For example, an adversary may have kept both private until two parallel honest proofs are generated. In this scenario, the red and black proofs with incremental height $3$ cannot be combined since they have different parents and the fork resolves later on according to the fork choice rule described in Section III-F.
@@ III-B Ledger Election
III-B1 Example proof model
@@ III-E Ledger Content
Further, as the transaction fees of the winning ledger of height h2h-2 are shared between the miner who created that ledger and miners of winning cluster at block height h1h-1, which are pointed by miners of height $h$, those fee payments have to be finalized in the ledger of height $h$ as well, which may require a smart contract algorithm involving the associated ledgers and the miners addresses of the previous two heights. The fraction of transaction fees that will be paid to the ledger creator is denoted as $r$ and will be explained in Section IV-C.
@@
III-F Difficulty Adjustment, Fork Choice and Block Confirmation Rules
@@
Note that, as the fork choice rule is the chain that contains the most aggregate work at all times, this implies that the ledger offering the most fees is only a temporary fork choice rule. More specifically, when a block with $L$ proofs is created and the ledger offering the most fees is shared, honest miners try to mine an initiator proof for the next block height referring to this ledger. However, as soon as the next initiator proof is created, this proof contains the most aggregate work and is the fork choice rule no matter which ledger it refers to as long as that ledger is semantically valid. We consider such a situation in more detail in Section D and show that it does not create any splits in honest views.
III-G Reward Distribution
The block at height $h$ receives total mining rewards proportional to the total proof height, $C(\omega^{h}_{L})$, which is distributed in the ledger of block at height $h+1$. To mitigate the effect of the selfish mining attacks, if there are two proofs which have the same proof height, i.e., $C(\omega^{h}_{i})=C(\omega^{h}_{j})$ for some miners $i$ and $j$, then they split the reward of the proof height between themselves. In other words, a proof $i$ for block height $h$ receives $\frac{1}{C(\omega^{h}_{L})}$ of the total block reward if it is the sole proof at that proof height, i.e., $C(\omega^{h}_{i})\neq C(\omega^{h}_{j})$ for any $j\neq i$, on the other hand if there is a parallel proof, each gets $\frac{1}{2C(\omega^{h}_{L})}$ of the total block reward.
@@ III-G Reward Distribution
III-G1 Example proof model
@@ III-H Adversarial Ledgers
To decrease the effect of adversarial ledger not being shared in a grace period which slightly delays the transition to the next height, we could allow a bit more than $L+1$ proofs, where the reward split for extra proofs could be omitted as well if the ledger offering most fees is not shared. We leave these minor adjustments to specific protocol implementations. Note that, as it is common in the literature and the practice, we assume honest miners echo the messages they receive, hence all honest miners eventually receive the adversarial ledger if one of them receives.
IV Analysis of the proposed protocol
@@
IV-A Consistency attacks
For the purpose of analyzing the double-spending probability, here we make same assumptions about the network model and the miners. Our assumptions resemble the state-of-the-art analysis of double-spending in bounded delay models for Nakamoto consensus [40, 31, 32, 41, 42]. As analysis of double-spending probability in the bounded delay models requires a good understanding of the possible adversarial attacks within the delay intervals, we advise the reader to check the cited papers. Here, we try to explain the analysis in as simple and concise terms as possible.
IV-A1 System Model
Assume that there are $n$ honest miners ($n$ is large) connected with an underlying network for communication, each holding an infinitesimal hashpower. Further, each proof generated by an honest miner experiences a network delay of $0,\delta]$, whereas each block ledger experiences a network delay of $0,\Delta]$. A single colluding adversarial miner is present in the system holding $\alpha$ fraction of the total hashpower. This adversarial entity experiences no communication delay and is aware of any proof as soon as it is generated. It can delay the transmission of the honest proofs/ledgers by maximally allowed $\delta$$\Delta$ for its own advantage and the ties are broken in its favor.
IV-A2 Confirmation Rule
Here, we only consider the double-spending probability for parameters $L\geq 10$, $b=1$ and give an upper bound for it. To repeat, our $b$-block confirmation rule states the following: A ledger together with the transactions inside are confirmed if it is the part of the heaviest chain and there are $b$-complete blocks mined on top of it. Notice that, in Nakamoto consensus, a $b$-block confirmation rule also counts the block that contains the ledger, therefore, when we compare our results, we should take this difference into account. More formally, our $b=1$ rule corresponds to $1\leq b\leq 2$ in Nakamoto consensus, since the ledger is proposed in one block and gets voted in the next (even though it remains hidden in the former block).
We refer the reader to Appendix E for the description of the double-spending attack and the related analysis which relies on the state-of-the-art techniques for Nakamoto consensus [40, 31, 32, 41, 42].

### v4

# An Incentive-Compatible Semi-Parallel Proof-of-Work Protocol

@@ Abstract
Parallel Proof-of-Work (PoW) protocols have been suggested in the literature to improve the safety guarantees, transaction throughput and confirmation latencies of Nakamoto consensus. In this work, we first consider the existing parallel PoW protocols and develop hard-coded incentive attack structures. Our theoretical results and simulations show that the existing parallel PoW protocols are more vulnerable to incentive attacks than the Nakamoto consensus, e.g., attacks have smaller profitability threshold and they result in higher relative rewards. Next, we introduce a voting-based semi-parallel PoW protocol that outperforms both Nakamoto consensus and the existing parallel PoW protocols from most practical perspectives such as communication overheads, throughput, transaction conflicts, incentive compatibility of the protocol as well as a fair distribution of transaction fees among the voters and the leaders. We use state-of-the-art analysis to evaluate the consistency of the protocol and consider Markov decision process (MDP) models to substantiate our claims about the resilience of our protocol against incentive attacks.
1 Introduction
@@
We introduce several rules for honest miners to make sure all miners of size $\alpha<0.5$ pick the proposal ledger that offers the most fees as $L$ increases. As a result, only one hidden ledger is shared and communication overheads as well as transaction conflicts are avoided. We introduce punishment mechanisms as in [20, 18, 19] to further reinforce incentive compatibility of the protocol as well as a fair distribution of transaction fees among the voters and the leaders. The introduction of ledger choice rule and split rule for the distribution of transaction fees makes sure that arbitrary manipulations of ledgers are not feasible for adversaries. We adapt the state-of-the-art security latency analysis of Nakamoto consensus [31, 32] to analyze the double-spending probability for our protocol. Further, we consider Markov decision process (MDP) models to substantiate our claims about the resilience of our protocol against incentive attacks.
In the rest of this paper, we assume a single colluding adversarial entity which holds $\alpha$ fraction of the hashpower in the system. We assume a network of $n$ honest miners connected to each other with some network topology and large $n$. We denote the network influence of the adversary, i.e., the fraction of honest miners that prefer the adversarial chain in case of a fork as $\gamma$. We abstract the mining puzzles with a collision-free hash function, $f_{H}$, which acts as a random oracle. In Section 2, we summarize the existing variations of the parallel PoW protocols and show their vulnerabilities against incentive attacks. In Section 3, we formally introduce our protocol and subsequently analyze it in Section 4. In Section 5, we discuss how our protocol relates to the existing work on PoW based protocols in general, its limitations and possible future work.
2 Incentive Attacks on Existing Variations of Parallel PoW Protocols
@@
2.1 Bobtail
In Bobtail, each proof has a hash value and can be created in parallel as usual and any proof with a hash value lower than a certain threshold is a valid proof. [^1] In the actual paper of [4], the average hash value of $L$ proofs needs to be below a threshold, however, for practical purposes and simplicity, we ignore proofs above a certain threshold and assume any proof below it is a valid proof as in [33]. The proof that has the lowest hash value is allowed to combine the current $L$ proofs to create a block. To prevent the proof withholding attacks, each proof has a support value, which is the main difference from the raw form. This support value is essentially a pointer to the lowest hash valued proof it has seen so far among the proofs for the current block height. To create a block, a lowest hash proof is only allowed to combine the proofs that have higher hash and support hash value.
@@ 2.1 Bobtail
To further strengthen the protocol, the authors consider a bonus reward rule where proofs, whose support proof is the lowest hash valued proof which eventually created the final block, get a bonus reward. The protocol also defines further incentive rules about prioritizing inclusion of proofs in case multiple sets of $L$ proofs are possible; see [4] for further details.
In Appendix A, we first provide details of the two attacks considered in the original paper of Bobtail, and later create a stronger hard-coded attack for which a pseudo-code is also given. We also provide both a theoretical overview as well as a simulation result which show Bobtail simply has profitability threshold of $\alpha_{T}=0$ due to the informational advantage that the adversary possesses. As a result, even a small home miner can mount an incentive attack on the protocol. It is also noted that the attack scales block interval times, decreases the throughput and increases the latency.
2.2 Tree- and DAG-Style Voting
Tailstorm [29] is a tree-style parallel PoW protocol [30]. Tailstorm introduces two main innovations compared to the raw form of the parallel PoW protocol. First, the proofs that are being generated for the current block height can refer to each other, creating a tree of proofs. As long as the root of the tree is the previous block, they are all eligible to be combined together to create the next block. Tailstorm further deviates from the raw form in how it rewards the proofs. Each block is rewarded proportional to the maximal depth of its proof tree and the reward is distributed equally to each proof in the tree. In Tailstorm, as every proof in a block gets rewarded equally, this also means that they are punished equally, since they can get more rewards if they build their proofs in a sequential chain and if a proof creates a fork, the reward for every proof inside the block decreases equally.
The addition of this tree structure and the new reward mechanism aim to mitigate the proof withholding attacks. DAG-style voting [30] introduces modifications along the same lines that aim to fix the unfair punishment mechanism in Tailstorm. In DAG-style voting, proofs are allowed to refer to multiple proofs which results in DAG. In this case, the reward of each proof is proportional to the total number of ancestor and descendant proofs it has within the block, which punishes the non-sequentiality more fairly; see [30] for further details.
@@ 2.2 Tree- and DAG-Style Voting
2.3 Numerical Results for the Incentive Attacks
@@
The handcrafted attacks we consider work irrespective of the $\gamma$ value, hence better attacks exist that take advantage of $\gamma$. Moreover, the handcrafted attacks are non-dynamic in the sense that the adversary does not change its strategy even when it encounters a high variance mining time. For example, in selfish mining attack of Nakamoto consensus, the profit of the attack comes from the fact that the adversary takes advantage of the high variance of the mining times, i.e., whenever the honest miners encounter a high block interarrival time, the adversary has an opportunity to be conversely lucky [4]. For example, since rewards are scaled according to the depth, the adversary could benefit from dynamically switching to honest mining for blocks for which it has created less than usual proofs.
3 Proposed Protocol Description
@@
Notice that, if a protocol separates the proofs from the ledgers, it decreases propagation times of proofs. As a result, less proofs will be mined during delay periods and thus the level of parallelness does not need to be high. This however, requires a ledger election mechanism that is incentive compatible. Further, in this case, ledger conflicts can be omitted with a clever mechanism. In this section, we explain a protocol that allows a reasonable amount of parallelness with small proof propagation times, an incentive compatible leader election mechanism, less communication overhead and has no ledger conflicts. Further, it has superior variance properties compared to Nakamoto consensus, i.e., a more fair distribution of coins, better safety guarantees and resilience against withholding attacks and throughput increase.
3.1 Blocks and Chain Structure
@@
A valid initiator proof consists of 6 parts concatenated together, i.e., $\omega^{h}_{1}=(\omega^{h}_{1,1},\ldots,\omega^{h}_{1,6})$ where w.l.o.g. we assume the miner who mines $\omega^{h}_{1}$ is the miner $i=1$. The sixth part of the proof, i.e., $\omega^{h}_{1,6}$, contains the nonce to make sure $f_{H}(\omega^{h}_{1})<T_{L}$. The fifth part, i.e., $\omega^{h}_{1,5}$, is simply the hash of the coin address of the miner $i=1$ to store its mining rewards in case its solution is accepted. The fourth part, $\omega^{h}_{1,4}$, is the Merkle root of the transaction ledger proposed by miner $1$, which will be explained in detail later. The third part, $\omega^{h}_{1,3}$, is simply the total accumulated transaction fees of the proposed ledger associated with $\omega^{h}_{1,4}$. The second part is a summary of (w.l.o.g. assume $L$) clustered proofs of preceding block height, i.e., $\omega^{h}_{1,2}=\omega^{h-1}=(f_{H}(\omega^{h-1}_{1}),\ldots,f_{H}(\omega^{h-1}_{L}))$. If two initiator proofs at height h1h-1 exist, say miner $1$ and miner $2$ created valid initiator proofs $\omega^{h}_{1}$ and $\omega^{h}_{2}$, then, they need to agree about the history to be clustered together in a block in order to prevent double-spending, i.e., $\omega^{h-1}_{1,2}=\omega^{h-1}_{2,2}$ ($\omega^{h-1}_{1,1}=\omega^{h-1}_{2,1}$ should also be satisfied, which will be explained in Section 3.2).
@@
3.1.1 Example Proof Model
@@
In a similar manner, at block height h1h-1, the proof represented with a red circle cannot be combined with the proof represented with a blue circle, since their parent sets differ. Regarding the red circles of block height $h+2$, multiple scenarios are possible. For example, an adversary may have kept both private until two parallel honest proofs are generated. In this scenario, the red and black proofs with incremental height $3$ cannot be combined since they have different parents and the fork resolves later on according to the fork choice rule described in Section 3.6.
@@ 3.2 Ledger Election
3.2.1 Example Proof Model
@@ 3.5 Ledger Content
Further, as the transaction fees of the winning ledger of height h2h-2 are shared between the miner who created that ledger and miners of winning cluster at block height h1h-1, which are pointed by miners of height $h$, those fee payments have to be finalized in the ledger of height $h$ as well, which may require a smart contract algorithm involving the associated ledgers and the miners’ addresses of the previous two heights. The fraction of transaction fees that will be paid to the ledger creator is denoted as $r$ and is explained in Section 4.3.
@@
3.6 Difficulty Adjustment, Fork Choice and Block Confirmation Rules
@@
Note that, as the fork choice rule is the chain that contains the most aggregate work at all times, this implies that the ledger offering the most fees is only a temporary fork choice rule. More specifically, when a block with $L$ proofs is created and the ledger offering the most fees is shared, honest miners try to mine an initiator proof for the next block height referring to this ledger. However, as soon as the next initiator proof is created, this proof contains the most aggregate work and is the fork choice rule no matter which ledger it refers to as long as that ledger is semantically valid. We consider such a situation in more detail in Appendix D and show that it does not create any splits in honest views.
3.7 Reward Distribution
The block at height $h$ receives total mining rewards proportional to the total proof height, $C(\omega^{h}_{L})$, which is distributed in the ledger of block at height $h+1$. To mitigate the effect of the selfish mining attacks, if there are two proofs which have the same proof height, i.e., $C(\omega^{h}_{i})=C(\omega^{h}_{j})$ for some miners $i$ and $j$, then they split the reward of the proof height among themselves. In other words, a proof $i$ for block height $h$ receives $\frac{1}{C(\omega^{h}_{L})}$ of the total block reward if it is the sole proof at that proof height, i.e., $C(\omega^{h}_{i})\neq C(\omega^{h}_{j})$ for any $j\neq i$, on the other hand if there is a parallel proof, each gets $\frac{1}{2C(\omega^{h}_{L})}$ of the total block reward.
@@ 3.7 Reward Distribution
3.7.1 Example Proof Model
@@ 3.8 Adversarial Ledgers
To decrease the effect of adversarial ledger not being shared in a grace period which slightly delays the transition to the next height, we could allow a bit more than $L+1$ proofs, where the reward split for extra proofs could be omitted as well if the ledger offering most fees is not shared. We leave these minor adjustments to specific protocol implementations. Note that, as it is common in the literature and the practice, we assume honest miners echo the messages they receive, hence all honest miners eventually receive the adversarial ledger if one of them receives it.
4 Analysis of the Proposed Protocol
@@
4.1 Consistency Attacks
For the purpose of analyzing the double-spending probability, here we make same assumptions about the network model and the miners. Our assumptions resemble the state-of-the-art analysis of double-spending in bounded delay models for Nakamoto consensus [40, 31, 32, 41, 42]. As analysis of double-spending probability in the bounded delay models requires a good understanding of the possible adversarial attacks within the delay intervals, we assume sufficient familiarity with the models/attacks/approaches in the cited blockchain analysis papers. Here, we try to explain the analysis in as simple and concise terms as possible. Note that, by showing that the upper bound on double-spending decreases exponentially, we essentially prove both the safety and the liveness properties under $\Delta$-synchronous settings due to the occurrence of Nakamoto blocks [40].
4.1.1 System Model
Assume that there are $n$ honest miners ($n$ is large) connected with an underlying network for communication, each holding an infinitesimal hashpower. Further, each proof generated by an honest miner experiences a network delay of $0,\delta]$, whereas each block ledger experiences a network delay of $0,\Delta]$. A single colluding adversarial miner is present in the system holding $\alpha$ fraction of the total hashpower. This adversarial entity experiences no communication delay and is aware of any proof as soon as it is generated. The adversary can delay the transmission of the honest proofs/ledgers by maximally allowed $\delta$$\Delta$ for its own advantage and the ties are broken in the favor of the adversary.
4.1.2 Confirmation Rule
Here, we only consider the double-spending probability for parameters $L\geq 10$, $b=1$ and give an upper bound for it. To recap, our $b$-block confirmation rule states the following: A ledger together with the transactions inside are confirmed if it is the part of the heaviest chain and there are $b$-complete blocks mined on top of it. Notice that, in Nakamoto consensus, a $b$-block confirmation rule also counts the block that contains the ledger, therefore, when we compare our results, we should take this difference into account. More formally, our $b=1$ rule corresponds to $1\leq b\leq 2$ in Nakamoto consensus, since the ledger is proposed in one block and gets voted in the next (even though it remains hidden in the former block).
We refer the reader to Appendix E for the description of the double-spending attack and the related analysis which relies on the state-of-the-art techniques for Nakamoto consensus; please see also [40, 31, 32, 41, 42].